OpenStudy (anonymous):

Indicate which property is illustrated in Step 6. Step 1 3 - 4x - 3 + 5x = (3 - 4x) + (-3 + 5x) Step 2 = (3 - 4x) + (5x - 3) Step 3 = 3 + (-4x + 5x) - 3 Step 4 = 3 + (-4 + 5)x - 3 Step 5 = 3 + x - 3 Step 6 = x + 3 - 3 Step 7 = x + 0 Step 8 = x @eyust707 @ganeshie8 @satellite73

OpenStudy (anonymous):

commutative law maybe?

OpenStudy (anonymous):

\[a+b=b+a\]so \[3+x-3=3-3+x\]

OpenStudy (anonymous):

Yeah, that makes the most sense I guess. :)